# Transformer-based RGB-T Tracking with Channel and Spatial Feature Fusion

### v1

@@ Abstract
Complementary RGB and TIR modalities enable RGB-T tracking to achieve competitive performance in challenging scenarios. Therefore, how to better fuse cross-modal features is the core issue of RGB-T tracking. Some previous methods either insufficiently fuse RGB and TIR features, or depend on intermediaries containing information from both modalities to achieve cross-modal information interaction. The former does not fully exploit the potential of using only RGB and TIR information of the template or search region for channel and spatial feature fusion, and the latter lacks direct interaction between the template and search area, which limits the model’s ability to fully exploit the original semantic information of both modalities. To alleviate these limitations, we explore how to improve the performance of a visual Transformer by using direct fusion of cross-modal channels and spatial features, and propose CSTNet. CSTNet uses ViT as a backbone and inserts cross-modal channel feature fusion modules (CFM) and cross-modal spatial feature fusion modules (SFM) for direct interaction between RGB and TIR features. The CFM performs parallel joint channel enhancement and joint multilevel spatial feature modeling of RGB and TIR features and sums the features, and then globally integrates the sum feature with the original features. The SFM uses cross-attention to model the spatial relationship of cross-modal features and then introduces a convolutional feedforward network for joint spatial and channel integration of multimodal features. Comprehensive experiments show that CSTNet achieves state-of-the-art performance on three public RGB-T tracking benchmarks. Code is available at https://github.com/LiYunfengLYF/CSTNet.
@@ I Introduction
Figure 1: Comparison between our cross-modal feature fusion method with other Transformer-based RGB-T trackers. (a) TIR features are fused into RGB features by extracting prompts, such as ViPT [1] (b) RGB and TIR features interact through the intermediate tokens, such as TBSI [2] (c) Our approach achieves cross-modal interaction through channel and spatial feature fusion of RGB and TIR features.
RGB-T tracking combines the advantages of the visual modality, which contains more information, and the infrared modality, which is less affected by the environment, to achieve stable tracking of the target in low light and adverse weather conditions. Due to its advantage in processing visual and infrared image sequences, it has a wide range of potential applications in traffic monitoring [3], robot perception [4] and autonomous driving [5].
As a multimodal vision task, how to achieve effective cross-modal interaction plays a crucial role in RGB-T tracking. In previous works, MDNet-based trackers propose a large number of cross-modal feature fusion methods, such as channel feature fusion [6], adapter [7], challenge-aware model [8], spatial-temporal feature fusion [9], multi-scale feature fusion [10], etc. These methods greatly advance the development of RGB-T tracking. However, the MDNet architecture itself makes it difficult for these trackers to extract discriminative deep feature. Then Siamese-based trackers [11][12][13] further explore the channel fusion methods of cross-modal deep features. However, their sample candidate boxes (RoIs) lack a global receptive field , as only local multimodal features of the search area are used for fusion. Overall, these MDNet-based and Siamese-based methods are limited by their own architecture and do not fully explore the potential of channel and spatial feature fusion. With the introduction of Transformer in RGB-T tracking, some methods use visual prompt learning [1][14] or adapter [15] to achieve better modeling of RGB feature guided by TIR feature, as illustrated in Figure 1 (a), while others are similar to the above methods, as illustrated in Figure 1 (b), like using feature bridging fusion [2], spatial-temporal feature fusion [16], etc. to achieve effective modeling of cross-modal features. Although these methods effectively achieve cross-modal feature interaction based on transformers, they are implemented by extracting intermediate prompt features or using partially fused cross-modal tokens. Their indirect feature interaction leads to a lack of use of the original semantics of multimodal features.
@@
As a basic feature fusion method, channel and feature fusion has significant advantages in the pixel-level interaction between RGB and TIR features. It directly models the original semantics of RGB and TIR features, thus avoiding the loss of key features caused by the use of intermediate tokens. For other RGB-T visual tasks, such as RGB-T semantic segmentation and RGB-T saliency detection, many cross-modal channel and spatial feature fusion modules have been proposed for Transformer-based methods, such as using multimodal channel attention to supplement key information between two modalities and using cross-attention to model spatial representations across-modalities [17], using feature concatenation, channel attention, and multi-level feature fusion module to recursively fuse intermediate features from different stages [18], etc. The success of the above feature fusion module demonstrates the enormous potential for direct interaction between RGB and TIR features in Transformer.
Therefore, this paper aims to combine channel and feature fusion with Visual Transformer (ViT). Specifically, we investigate how to achieve direct interaction of multimodal features in ViT and improve the performance of our tracking pipeline by using the original semantic information from the cross-modal features. Most transformer-based methods design an independent feature fusion module to preserve the powerful feature extraction and relationship modeling capabilities of the ViT. Following this design, the critical problem becomes how to design a channel feature fusion module and a spatial feature fusion module.
In this work, we propose an RGB-T tracker referred to as CSTNet. CSTNet explores the advantages of directly integrating RGB and TIR features in space and channels in ViT. Similar to [2], we design a cross-modal channel feature fusion module (CFM) and a spatial feature fusion (SFM) module, respectively. We cascade them into a block, and insert the blocks into the ViT, which is our backbone. The CFM performs joint channel enhancement and multi-level spatial modeling of multimodal features, and then globally integrates the original semantic and fused semantic information across modal features. The SFM achieves cross-modal feature interaction through cross-attention, and then combines spatial and channel features through convolutional feedforward layers.
@@
(2) We propose a cross-modal spatial feature fusion module (SFM) that first aggregates local spatial feature, then models the spatial relationships of multimodal features using cross-attention module, and finally integrates the spatial and channel information of multimodal features through a convolutional feedforward network.
@@ II Related Work
II-A RGB-T Tracking
Some RGB-T trackers propose channel fusion and spatial fusion methods for cross-modal features. For example, DAFNet [6] proposes a channel-attention-based recursive fusion chain to achieve cross-modal and multi-level feature fusion in MDNet. SiamCDA [11] proposes a complementarity aware feature fusion module. It concatenates RGB and TIR features, and then enhances cross-modal features through a joint channel attention module. Finally, it achieves feature fusion through cross-modal residual connection and feature concatenation. SiamTDR [12] concatenates RGB and TIR features and then calculates the weight of channel attention by dynamic convolution to obtain fused feature. The fusion module of SiamMALL [13] is similar to the above methods, except that it replaces the feature concatenation of the last layer with feature addition. MACFT [19] proposes a cross-attention fusion module and a mix-attention fusion module for spatial feature fusion in Transformer. It only performs cross-attention operations on the multimodal features output by its backbone to obtain fused features.
Compared with these channel feature fusion modules, our method not only uses channel enhancement, but also introduces multi-level local spatial feature modeling and global integration module to achieve cross-modal feature fusion. Compared to the spatial fusion module proposed by MACFT [19], our method explores the potential of cross-attention and integrates the spatial representation of the fused feature by convolutional feedforward network. In addition, we insert our modules into different layers of the backbone to achieve cross-modal feature interactions at different stages.
In addition, there are many other effective cross-modal feature fusion methods in RGB-T tracking. BAT [15] proposes a universal bi-directional adapter to achieve interaction between RGB and TIR features in transformer blocks. TBSI [2] uses the fused template as a bridge to achieve information interaction between RGB and TIR modalities in the search area based on the attention mechanism. CAT++ [20] proposes five different attribute-based branches to perceive different challenges, and then fuses the features of different branches through an aggregation interaction module to achieve multimodal tracking. TATrack [16] introduces an online template to model temporal information, and achieves cross-modal feature interaction between the initial template and the online template.
@@ II-A RGB-T Tracking
II-B Feature fusion for other RGB-T visual tasks
@@
Compared to these methods, our method uses the fused feature to help model the association between the template and the search area, rather than using the fused feature to distinguish the category of pixels or to predict the target’s boxes. As a result, our modules differ significantly in terms of insertion layers and model design.
@@ III Method
The overall framework of CSTNet is shown in Figure 2. We first embed the template images and search area images into tokens, and then fed the tokens into transformer blocks. Our proposed channel feature fusion module (CFM) and spatial feature fusion module (SFM) are inserted between transformer blocks for cross-modal feature interaction. Finally, the RGB and TIR features of the search area output from the last transformer block are added and fed into the prediction head to obtain the target state. Due to the same processing of template and search area in CFM and SFM, we only show the processing of multimodal features in the search area in sections III-B and III-C.
@@ III-A Backbone
We follow the recent RGB-T trackers [2][16] to select ViT [22] as our backbone. The inputs of the backbone are the RGB and TIR template images represented as $z_{r},z_{t}\in R^{H_{z}\times W_{z}\times 3}$ and the RGB and TIR search area images are represented as $x_{r},x_{t}\in R^{H_{x}\times W_{x}\times 3}$. These four images are first embedded as tokens and positional embeds are added as follows:
@@
where PE denotes Patch Embed function as same in [22]. $Z_{r},Z_{t}\in R^{N_{z}\times C}$ denote the init feature of RGB and TIR template tokens. $N_{z}=H_{z}W_{z}/p^{2}$, where $p$ is the down-sampling stride of patch embedding. $C$ is the channel number. $X_{r},X_{t}\in R^{N_{x}\times C}$ denote the init feature of RGB and TIR seach area tokens, where $N_{z}=H_{z}W_{z}/p^{2}$. $E_{z}$ and $E_{x}$ are the positional embeds with the same shape as $Z_{r}$ and $X_{r}$, respectively.
Since the feature extraction and modeling is the same for both modalities, we only discuss the process in the RGB modality. The RGB tokens are concatenated as $H_{r}=[X_{r};Z_{r}]\in R^{(N_{x}+N_{z})\times C}$, then fed $H_{r}$ into a series of Transformer [22] blocks.
@@
$\begin{split}A&=\text{Softmax}(\frac{QK^{T}}{\sqrt{C}})=\text{Softmax}(\frac{[X_{qrt};Z_{qrt}][X_{krt};Z_{krt}]^{T}}{\sqrt{C}})\\ &=\text{Softmax}(\frac{[X_{qrt}X_{krt}^{T},X_{qrt}Z_{krt}^{T};Z_{qrt}X_{krt}^{T},Z_{qrt}Z_{krt}^{T}]}{\sqrt{C}})\end{split}$ (2)
@@ III-B Channel feature fusion module
The local spatial aggregation module (LSA) is used to extract multi-level local spatial features and model key representations of mixed multi-level features through a nonlinear layer. Specifically, a linear transformation is first performed on the feature $X_{c}$ and obtain $X_{c}^{\text{fc}}=\text{Linear}(X_{c})\in R^{N_{x}\times C}$.
@@
Another linear layer is used to perform a linear transformation on the fused features, integrating multi-level features and improving its consistency, represented as $X_{c}^{\text{lsa}}=\text{Linear}(\text{Gelu}(\text{Conv}_{1\times 1}(X_{c}^{\text{msf}})))\in R^{N_{x}\times C}$, where Gelu represents GELU activate function [24].
After parallel SE and LSA modules, the features are added, represented as $X_{c}^{\text{add}}=X_{c}^{\text{se}}+X_{c}^{\text{lsa}}\in R^{N_{x}\times C}$. In addition, residual connections are introduced to supplement the original semantic information lost in multimodal features, represented as $X_{r}^{\text{res}}=X_{r}+X_{c}^{\text{add}}$ and $X_{t}^{\text{res}}=X_{t}+X_{c}^{\text{add}}$.
The global integration module (GIM) concatenates RGB and TIR features, and fed it into a multi-layer perceptron (MLP) with residual connection. Then, split the concatenated feature to obtain integrated RGB and TIR features. The process is described as:
@@
where Split maps the feature from $R^{N_{x}\times 2C}$ to two $R^{N_{x}\times C}$ features. MLP is consist of two linear layers. The $X_{r}^{\text{cfm}}$ and $X_{t}^{\text{cfm}}$ are the CFM output features.
Due to the difference between the template and the search area, GIM is not shared when processing the features between the template and the search area.

### v2

@@ Abstract
Complementary RGB and TIR modalities enable RGB-T tracking to achieve competitive performance in challenging scenarios. Therefore, how to better fuse cross-modal features is the core issue of RGB-T tracking. Some previous methods either insufficiently fuse RGB and TIR features, or depend on intermediaries containing information from both modalities to achieve cross-modal information interaction. The former does not fully exploit the potential of using only RGB and TIR information of the template or search region for channel and spatial feature fusion, and the latter lacks direct interaction between the template and search area, which limits the model’s ability to fully exploit the original semantic information of both modalities. To alleviate these limitations, we explore how to improve the performance of a visual Transformer by using direct fusion of cross-modal channels and spatial features, and propose CSTNet. CSTNet uses ViT as a backbone and inserts cross-modal channel feature fusion modules (CFM) and cross-modal spatial feature fusion modules (SFM) for direct interaction between RGB and TIR features. The CFM performs parallel joint channel enhancement and joint multilevel spatial feature modeling of RGB and TIR features and sums the features, and then globally integrates the sum feature with the original features. The SFM uses cross-attention to model the spatial relationship of cross-modal features and then introduces a convolutional feedforward network for joint spatial and channel integration of multimodal features. Comprehensive experiments show that CSTNet achieves state-of-the-art performance on three public RGB-T tracking benchmarks. To promote the practicality of the model, we retrain the model with CSNet as the pre-training weights in the model with CFM and SFM removed, and propose CSTNet-small, which achieves 36% reduction in parameters and 24% reduction in Flops, and 50% speedup with a 1-2% performance decrease. Code is available at https://github.com/LiYunfengLYF/CSTNet.
@@ I Introduction
Figure 1: Comparison between our cross-modal feature fusion method with other Transformer-based RGB-T trackers. (a) TIR features are fused into RGB features by extracting prompts, such as ViPT [1] (b) RGB and TIR features interact through the intermediate tokens, such as TBSI [2] (c) Our method achieves cross-modal interaction through channel and spatial feature fusion of RGB and TIR features.
RGB-T tracking is a combination of single object tracking (SOT) and multi-modal vision task that provides multi-modal appearance templates of the target in the initial frame and then predicts the position and scale of the target in the multi-modal images in subsequent frames. It combines the advantages of the visual modality, which contains more information, and the infrared modality, which is less affected by the environment, to achieve stable target tracking in low light and adverse weather conditions. Due to its advantage in processing visual and infrared image sequences, it has a wide range of potential applications in traffic monitoring [3], robot perception [4] and autonomous driving [5].
As a multi-modal vision task, how to achieve effective cross-modal interaction plays a crucial role in RGB-T tracking. It is worth noting that while some trackers [6] discuss the impact of RGB-T image misalignment on tracking performance, most current trackers are designed with the assumption of spatial registration of RGB-T images. In previous works, MDNet-based trackers have proposed a large number of cross-modal feature fusion methods, such as channel feature fusion [7], adapter [8], challenge-aware model [9], spatio-temporal feature fusion [10], multi-scale feature fusion [11], etc. These methods greatly advance the development of RGB-T tracking. However, the MDNet architecture itself makes it difficult for these trackers to extract discriminative deep features. Then, Siamese-based trackers [12][13][14] further explore the channel fusion methods of cross-modal deep features. However, their sample candidate boxes (RoIs) lack a global receptive field , because only local multimodal features of the search area are used for fusion. Overall, these MDNet-based and Siamese-based methods are limited by their own architecture and do not fully explore the potential of channel and spatial feature fusion. With the introduction of Transformer in RGB-T tracking, some methods use visual prompt learning [1][15] or adapter [16] to achieve better modeling of RGB feature guided by TIR feature, as illustrated in Figure 1 (a), while others are similar to the above methods, as illustrated in Figure 1 (b), like using feature bridging fusion [2], spatio-temporal feature fusion [17], etc. to achieve effectively model of cross-modal features. Although these methods effectively achieve cross-modal feature interaction based on transformers, they are implemented by extracting intermediate prompt features or using partially fused cross-modal tokens. Their indirect feature interaction leads to a lack of use of the original semantics of multi-modal features.
@@
As a basic feature fusion method, channel and feature fusion has significant advantages in the pixel-level interaction between RGB and TIR features. It directly models the original semantics of RGB and TIR features, thus avoiding the loss of key features caused by the use of intermediate tokens. For other RGB-T visual tasks, such as RGB-T semantic segmentation and RGB-T saliency detection, many cross-modal channel and spatial feature fusion modules have been proposed for Transformer-based methods, such as using cross-attention to model cross-modal feature representations [18], using channel attention and spatial feature fusion module to fuse intermediate features from different stages [19], etc. The success of the above feature fusion module demonstrates the enormous potential of direct interaction between RGB and TIR features in Transformer.
Therefore, this paper aims to combine channel and feature fusion with Visual Transformer (ViT). Specifically, we investigate how to achieve direct interaction of multimodal features in ViT and improve the performance of our tracking pipeline by using the original semantic information of the cross-modal features. Most transformer-based methods design an independent feature fusion module to preserve the powerful feature extraction and relationship modeling capabilities of ViT. Following this design, the critical problem becomes how to design a channel feature fusion module and a spatial feature fusion module.
In this work, we propose an RGB-T tracker referred to as CSTNet. CSTNet explores the advantages of directly integrating RGB and TIR features in space and channels in ViT. To the best of our knowledge, our work is the first to introduce direct interactions in ViT-based RGB-T trackers. Similar to [2], we design a cross-modal channel feature fusion module (CFM) and a spatial feature fusion (SFM) module. We cascade them into a block and insert the blocks into the ViT, which is our backbone. The CFM performs joint channel enhancement and multi-level spatial modeling of multimodal features, and then globally integrates the original semantic and fused semantic information across modal features. The SFM achieves cross-modal feature interaction through cross-attention, and then combines spatial and channel features through convolutional feedforward layers.
In addition, to promote the practicality of the model, we also propose a smaller version of CSTNet-small. It removes the CFM and SFM from the network and uses the CSTNet model as the pre-training weights. The CSTNet-small with fewer parameters and faster speed will promote the applications of our work in intelligent vehicles such as unmanned aerial vehicle (UAV), unmanned surface vehicle (USV), underwater vehicle (UV), etc.
@@
(2) We propose a cross-modal spatial feature fusion module (SFM) that first aggregates local spatial features, then models the spatial relationships of multimodal features using a cross-attention module, and finally integrates the spatial and channel information of multimodal features through a convolutional feedforward network.
@@
(4) We propose a smaller version CSTNet-small. Compared to CSTNet, it removes the CFM and SFM modules and uses CSTNet as the pre-training model and achieves over 36% reduction in parameters, over 24% reduction in Flops, and 50% speedup with 1-2% performance decrease.
@@ II Related Work
II-A RGB-T tracker based on Indirect Feature Interaction
Some RGB-T trackers propose channel fusion and spatial fusion methods for cross-modal features. DAFNet [7] proposes a channel-attention-based recursive fusion chain to achieve cross-modal and multi-level feature fusion in MDNet. SiamCDA [12] proposes a complementary aware feature fusion module. It concatenates RGB and TIR features, and then enhances cross-modal features through a joint channel attention module. Finally, it achieves feature fusion through cross-modal residual connection and feature concatenation. SiamTDR [13] concatenates RGB and TIR features and then calculates the weight of channel attention by dynamic convolution to obtain fused feature. The fusion module of SiamMALL [14] is similar to the above methods, except that it replaces the feature concatenation of the last layer with feature addition. MACFT [20] proposes a cross-attention fusion module and a mix-attention fusion module for spatial feature fusion in Transformer. It only performs cross-attention operations on the multimodal features output by its backbone to obtain fused features.
Compared with these channel feature fusion modules, our method not only uses channel enhancement, but also introduces multi-level local spatial feature modeling and global integration module to achieve cross-modal feature fusion. Compared to the spatial fusion module, our method explores the potential of cross-attention and integrates the spatial representation of the fused feature by convolutional feedforward network. In addition, we insert our modules into different layers of the backbone to achieve cross-modal feature interactions at different stages. Compared to these trackers, our method achieves continuous interaction of cross modal features in channels and space, while these methods only explore channel fusion or spatial feature interaction separately.
II-B RGB-T tracker based on Indirect Feature Interaction
BAT [16] proposes a universal bi-directional adapter to achieve interaction between RGB and TIR features in transformer blocks. TBSI [2] uses the fused template as a bridge to achieve information interaction between RGB and TIR modalities in the search area based on the attention mechanism. CAT++ [21] proposes five different attribute-based branches to perceive different challenges, and then fuses the features of different branches through an aggregation interaction module to achieve multimodal tracking. TATrack [17] introduces an online template to model temporal information, and achieves cross-modal feature interaction between the initial template and the online template.
@@ II-B RGB-T tracker based on Indirect Feature Interaction
II-C Feature fusion for other RGB-T visual tasks
@@
Compared to these methods, our method is different in terms of task form. Compared to the cross-attention module in CMX [18], our cross-attention-like module improves cross-modal feature interaction by aggregating local space features, sharing integration layers, maintaining cross-modal feature channel consistency during feature fusion, and decoupling the learning of target space and channel features.
@@ III Method
Our method is based on the assumption of spatial alignment of RGB-T images. The overall framework of CSTNet is shown in Figure 2. We first embed the template images and search area images into tokens, and then fed the tokens into transformer blocks. Our proposed channel feature fusion module (CFM) and spatial feature fusion module (SFM) are inserted between the transformer blocks for cross-modal feature interaction. Finally, the RGB and TIR features of the search area output from the last transformer block are added and fed into the prediction head to obtain the target state. Due to the identical processing of template and search area in CFM and SFM, we only show the processing of multimodal features in the search area in sections III-B and III-C. CSTNet-small removes the CFM and SFM modules from CSTNet and trains the modal with CSTNet weights as a pre-trained model.
@@ III-A Backbone
We follow the recent RGB-T trackers [2][17] to select ViT [23] as our backbone. The inputs of the backbone are the RGB and TIR template images, represented as $z_{r},z_{t}\in R^{H_{z}\times W_{z}\times 3}$ and the RGB and TIR search area images are represented as $x_{r},x_{t}\in R^{H_{x}\times W_{x}\times 3}$. These four images are first embedded as tokens, and positional embeds are added as follows:
@@
where PE is the Patch Embed function as in [23]. $Z_{r},Z_{t}\in R^{N_{z}\times C}$ denotes the init feature of RGB and TIR template tokens. $N_{z}=H_{z}W_{z}/p^{2}$, where $p$ is the down-sampling stride of the patch embedding. $C$ is the channel number. $X_{r},X_{t}\in R^{N_{x}\times C}$ denote the init feature of RGB and TIR seach area tokens, where $N_{x}=H_{x}W_{x}/p^{2}$. $E_{z}$ and $E_{x}$ are the position embeds with the same shape as $Z_{r}$ and $X_{r}$, respectively.
Since the feature extraction and modeling is the same for both modalities, we only discuss the process in the RGB modality. The RGB tokens are concatenated as $H_{r}=[X_{r};Z_{r}]\in R^{(N_{x}+N_{z})\times C}$, then $H_{r}$ is fed into a series of Transformer [23] blocks.
@@
$\small\begin{split}A&=\text{Softmax}(\frac{QK^{T}}{\sqrt{C}})V=\text{Softmax}(\frac{[X_{qrt};Z_{qrt}][X_{krt};Z_{krt}]^{T}}{\sqrt{C}})V\\ &=\text{Softmax}(\frac{[X_{qrt}X_{krt}^{T},X_{qrt}Z_{krt}^{T};Z_{qrt}X_{krt}^{T},Z_{qrt}Z_{krt}^{T}]}{\sqrt{C}})V\end{split}$ (2)
@@ III-B Channel feature fusion module
The local spatial aggregation module (LSA) is used to extract multi-level local spatial features and model key representations of mixed multi-level features by a non-linear layer. Specifically, a linear transformation is first performed on the feature $X_{c}$ to obtain $X_{c}^{\text{fc}}=\text{Linear}(X_{c})\in R^{N_{x}\times C}$.
@@
Another linear layer is used to perform a linear transformation on the fused features to integrate multi-level features and improve their consistency, represented as $X_{c}^{\text{lsa}}=\text{Linear}(\text{Gelu}(\text{Conv}_{1\times 1}(X_{c}^{\text{msf}})))\in R^{N_{x}\times C}$, where Gelu represents the GELU activate function [25].
After the parallel SE and LSA modules, the features are added, represented as $X_{c}^{\text{add}}=X_{c}^{\text{se}}+X_{c}^{\text{lsa}}\in R^{N_{x}\times C}$. In addition, residual connections are introduced to complement the original semantic information lost in multimodal features, represented as $X_{r}^{\text{res}}=X_{r}+X_{c}^{\text{add}}$ and $X_{t}^{\text{res}}=X_{t}+X_{c}^{\text{add}}$.
The global integration module (GIM) concatenates the RGB and TIR features and feeds it into a multi-layer perceptron (MLP) with residual connection. Then, the concatenated feature is split to obtain integrated RGB and TIR features. The process is described as:
@@
where Split maps the feature from $R^{N_{x}\times 2C}$ to two $R^{N_{x}\times C}$ features. MLP consists of two linear layers. The $X_{r}^{\text{cfm}}$ and $X_{t}^{\text{cfm}}$ are the CFM output features.
Due to the difference between the template and the search area, GIM is not shared between the template and the search pane when processing features.
III-C Spatial feature fusion module
Our proposed SFM module aims to model spatial relationships in cross-modal features and to integrate the channel and spatial features of the RGB and TIR modalities through a convolutional feedforward network. As shown in Figure 4, the SFM contains a local perception unit (LPU), a cross-attention module, and a convolutional feedforward network (CFN).